# Iterated Invariant Extended Kalman Filter (IterIEKF)

### v4

@@ I Introduction
In the field of state estimation and observer design, the Extended Kalman Filter (EKF) is one of the most widespread methods used in practice. However, it is based on linearization, and the associated errors were early recognized as possibly degrading performance [1]. This motivated the development of iterative filtering algorithms such as the iterated EKF (IterEKF[2, 3, 4], iterated EKF on Lie groups (LG-IterEKF[5, 6, 7], and smoothers [8], where the discrepancy between the nonlinear output function and its first-order approximation is reduced, by refining the operating point.
@@
On another note, observer design has benefited from geometric approaches over the past two decades. They may be traced back to attitude and pose estimation [9, 10, 11]. When turning to the more general problem of inertial navigation and localization, the Invariant Extended Kalman Filter (IEKF[12, 13] has become a key alternative to the standard EKF. Its theoretical properties include convergence guarantees [12], consistency properties in the presence of unobservability [13], see also related works [14, 15, 16], and have led to applications in various fields, e.g.,  [13, 8, 17, 18, 19, 20, 14] and in the industry [13]. The field of observers has benefited from the introduction of new groups brought by the IEKF theory, see e.g., [21, 22, 23], namely the groups $SE_{2}(3)$ and $SE_{k}(d)$ introduced in [24, 12] and their relation to the navigation equations, see [25] for a recent perspective. The equivariant observer framework, see [26], is closely related.
@@
In this technical note, we introduce the Iterated Invariant Extended Kalman Filter (IterIEKF), a refinement of the IEKF [13] that leverages the Gauss-Newton (GN) method to enhance its measurement update step, and which systematically improves its accuracy. Our main contribution is then a comprehensive analysis of the IterIEKF properties, revealing rare properties in nonlinear estimation that remind of the linear case, and offering new tools and insights into the theory of invariant filtering. Finally, to highlight the practical relevance of our approach, we apply the IterIEKF to a problem of engineering interest, and show it outperforms the IEKF when observations are very accurate (low measurement noise).
@@
While the ultimate goal of invariant filtering is to recover properties of the linear Kalman filter in a nonlinear setting, and to identify systems of interest that lend themselves to this goal, a parallel line of work has been concerned with how to accommodate in general the geometric nature of the state space, without seeking further properties. In particular, the “Lie-Group EKF” (LG-EKF) from [27] proposes a generic intrinsic version of the EKF on Lie groups, that accounts for geometric features such as curvature. An iterated version, called LG-IterEKF was proposed in [5]. See, e.g., [28] and applications to inertial-lidar navigation [29] for recent references in this field. When tailored to the specific class of problems we consider, namely by using the groups from invariant filtering and the choice of errors that it advocates, the LG-IterEKF (that we call then adapted) is close to the proposed IterIEKF, although slightly different. As a result, we can prove the properties of the IterIEKF carry over to this adapted LG-IterEKF. This novel result constitutes a secondary contribution.
@@
Section II introduces the proposed IterIEKF algorithm for both noisy and noise-free measurements. Section III derives theoretical properties of the IterIEKF in the context of noise-free measurements. We prove they carry over to the LG-IterEKF from [5] under some conditions. Section IV evaluates in simulations the performance of the proposed algorithm compared to other state-of-the-art filters for estimating the extended pose (orientation, velocity, and position) of a crane hook, equipped with an Inertial Measurement Unit (IMU).
@@ II-A The IEKF equations
The invariant framework assumes the state $\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}$ is an element of a matrix Lie group $G\subset GL_{N}(\mathbb{R})$ having dimension $n$, where $GL_{N}(\mathbb{R})$ denotes the group of $N\times N$ invertible matrices. An example of such a state $\boldsymbol{\chi}$ is given below, in (15). The exponential map of $G$ is defined as $\exp_{G}(\cdot):=\exp_{m}(\mathcal{L}_{\mathfrak{g}}(\cdot))$, where $\exp_{m}(\cdot)$ is the matrix exponential and $\mathcal{L}_{\mathfrak{g}}(\cdot)$ is the bijective linear map identifying the Lie algebra $\mathfrak{g}$ with $\mathbb{R}^{n}$, see e.g., [12]. The IEKF comes in two versions, the left- or the right-invariant version, depending on the form of the observations, see [13]. This note focuses on the left-IEKF. Transposition to the right-IEKF is straightforward.
@@
The (left) IEKF assumes that the state follows a concentrated Gaussian distribution on $G$ [30, 31, 32, 33]. At time $k$, letting $l$ denote current time index $k$, or previous index k1k-1,
@@
When the function $\mathbf{f}$ possesses the group affine property (i.e., is of the form $\mathbf{f}(\raisebox{1.0pt}{$\boldsymbol{\chi}$},\mathbf{u},\mathbf{w})=\bar{\mathbf{f}}(\raisebox{1.0pt}{$\boldsymbol{\chi}$},\mathbf{u})\mathbf{g}(\mathbf{w})$ where $\bar{\mathbf{f}}$ satisfies $\bar{\mathbf{f}}(\mathbf{a}\mathbf{b},\mathbf{u})=\mathbf{a}\bar{\mathbf{f}}(\mathbf{b},\mathbf{u})+\bar{\mathbf{f}}(\mathbf{a},\mathbf{u})\mathbf{b}-\mathbf{a}\bar{\mathbf{f}}(\mathbf{I},\mathbf{u})\mathbf{b}$, for all $\mathbf{a},\mathbf{b}\in G$, $\mathbf{u}\in\mathbb{R}^{m}$, see [13]), the Jacobian $\mathbf{F}_{k}$ becomes independent of the current state estimate $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}$ [12, 13]. Moreover, in the absence of process noise, this results in the exact error dynamics $\boldsymbol{\xi}_{k+1\mid k}=\mathbf{F}_{k}\boldsymbol{\xi}_{k\mid k}$, where the Jacobian $\mathbf{F}_{k}$, i.e., the first order, fully captures the nonlinearity of $\mathbf{f}$. This property called log-linearity is a key feature of invariant filtering. Note that an IEKF can be devised even if $\mathbf{f}$ is not group affine, based on expansion (4), and the theory to follow applies.
@@ II-B The IterIEKF equations
The optimization problem (6) does not admit a closed-form solution. The IEKF addresses this by approximating $\exp_{G}(\boldsymbol{\xi})$ using its first-order Taylor expansion around $\mathbf{0}$. Inspired by the iterated EKF [2], a more accurate solution can be obtained by applying the GN algorithm to iteratively refine the estimate until convergence. This process results in a sequence of updates that closely resemble those of the IEKF. The resulting Iterated IEKF (IterIEKF) is outlined in Algorithm 1. Details are provided in Appendix A-A.
@@
$\mathbf{\mathcal{J}}_{r}(\boldsymbol{\xi})$ denotes the right Jacobian [30] of $G$ at $\boldsymbol{\xi}$, see (30).
Iterating proves particularly useful when measurement noise is low, as we will see in the sequel. However, in the limit case where the measurement noise magnitude approaches zero, the innovation covariance $\mathbf{S}_{k}$ may become rank-deficient [34], in which case the Kalman gain is undefined. An adapted version of the IterIEKF for this specific scenario is detailed in Algorithm 2. Details are provided in Appendix A-B.
@@ III Theoretical properties of the IterIEKF
Deriving properties in the presence of measurement noise seems out of reach, as is often the case in nonlinear filtering (e.g., the log-linear error property of [12] only holds in the absence of process noise). However, we show that the IterIEKF possesses strong properties in the absence of measurement noise (whereas the state estimate and the propagation may be noisy). We therefore consider noise-free measurement
@@ III-A Properties of considered noise-free measurements
The theory of (left) invariant filtering focuses on measurements of the form (1b), see [12], which boil down to (7) when the noise is turned off. It turns out that the corresponding “observed set” (8) possesses an interesting structure.
@@
The first point of Proposition 2 is directly useful for proving the compatibility properties of the IterIEKF, in the sense of Definition 2. This will be established in the next subsection, relying on the following key Lemma 1. We also note, in passing, that the second point will only become relevant when extending the present theory to the iterated EKF on Lie groups (LG-IterEKF) of [33], under certain conditions.
@@ III-C Consequences of the compatibility property
Let us first consider the first point of Definition 2. During the update with $\mathbf{y}_{k}$, we see at line 18 of Algorithm 1 that $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}=\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k-1}\exp_{G}(\boldsymbol{\xi}_{k\mid k-1}^{i})$, where $\boldsymbol{\xi}_{k\mid k-1}^{i}=\mathbf{K}_{k}^{i}\mathbf{z}_{k}^{i}\in\operatorname{Im}\mathbf{K}_{k}^{i}\subseteq\operatorname{Im}\mathbf{P}_{k\mid k-1}$, with $\operatorname{Im}\mathbf{K}_{k}^{i}$ denoting the span of $\mathbf{K}_{k}^{i}$. Since $\bar{\mathbf{H}}_{k}\mathbf{P}_{k\mid k-1}\bar{\mathbf{H}}_{k}^{T}=\mathbf{0}$ by assumption, we necessarily have $\bar{\mathbf{H}}_{k}\boldsymbol{\xi}_{k\mid k-1}^{i}=\mathbf{0}$. Then, as $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k-1}\in\mathcal{S}_{\{\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}\bar{\mathbf{d}}_{k}=\bar{\mathbf{y}}_{k}\}}$, a direct application of Lemma 1 ensures $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}\in\mathcal{S}_{\{\raisebox{1.0pt}{$\boldsymbol{\chi}$}\bar{\mathbf{d}}_{k}=\bar{\mathbf{y}}_{k}\}}$.
@@ III-D Extending the theory to the iterated EKF on Lie groups
The LG-IterEKF, introduced in [5], transposes the standard iterated EKF to Lie group state spaces, by leveraging an intrinsic GN method on the group, that accounts for its manifold structure, notably curvature. LG-IterEKF of [5] is a general algorithm, meant to address general problems on Lie groups having no specific structure, beyond the intrinsic nature of the state space. By contrast, our IterIEKF addresses specific types of observations, namely (1b) (or its right-invariant counterpart $\mathbf{y}_{k}=\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}^{-1}\mathbf{d}_{k}+\mathbf{n}_{k}$, in which case an iterated right-invariant EKF should be used, by mimicking the difference between the LIEKF and the RIEKF of [12]). It defines a modified innovation $\mathbf{z}_{k}=\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k}^{-1}\mathbf{y}_{k}-\mathbf{d}_{k}$, allowing for the use of the invariant Kalman filter framework of [12, 13]. The standard invariant EKF has had various successes across control, navigation, and robotics, and is clearly a different algorithm than the LG-EKF of [27].
That said, the questions one could address are as follows. Assume we use the theory of invariant filtering to endow the LG-IterEKF with all the features of the IterIEKF—namely, measurements of the form (7), uncertainty model (2), and the Kalman gain from Algorithm 2 to address rank deficiency issues, as well as the appropriate choice regarding left or right versions (note that [5] chooses a systematic right-invariant error based setting whereas we advocate to choose the error depending on the form of the observations), resulting in an “invariant” version of the LG-IterEKF, that we call adapted LG-IterEKF, see Algorithm 3. Then, 1) do we actually recover the IterIEKF, and 2) if no, does it inherit any of the properties we have proved, given that no such properties have ever been proved for the LG-IterEKF? The answer to point 1 is no, as intrinsic GN on the group entails adding various Jacobians and differs from the GN algorithm addressing (6), that directly optimizes in a linear space. The answer to point 2 is yes, the properties above do carry over. This is detailed, proved, and discussed, in Appendix B.
Algorithm 3 The adapted LG-IterEKF from [5] when measurements (7) are considered, uncertainty model (2) is used, potential rank deficiency issues in the Kalman gain are addressed via the noise-free gain from Algorithm 2, and a left-invariant error is used instead. See (30) for definition of $\mathbf{\mathcal{J}}_{r}$.

### v5

@@ I Introduction
In the field of state estimation and observer design, the Extended Kalman Filter (EKF) is one of the most widespread methods used in practice. However, it is based on linearization, and the associated errors were early recognized as possibly degrading performance [schmidt1981kalman]. This motivated the development of iterative filtering algorithms such as the iterated EKF (IterEKF[bell1993iterated, zhao2016robust, bloesch2017iterated], iterated EKF on Lie groups (LG-IterEKF[bourmaud2016intrinsic, liu2023general, chahbazian2021laplace], and smoothers [van2020invariant], where the discrepancy between the nonlinear output function and its first-order approximation is reduced, by refining the operating point.
@@
On another note, observer design has benefited from geometric approaches over the past two decades. They may be traced back to attitude and pose estimation [bonnabel2005invariant, mahony2008nonlinear, cohen2020navigation]. When turning to the more general problem of inertial navigation and localization, the Invariant Extended Kalman Filter (IEKF[barrau2016invariant, barrau2018invariant] has become a key alternative to the standard EKF. Its theoretical properties include convergence guarantees [barrau2016invariant], consistency properties in the presence of unobservability [barrau2018invariant], see also related works [mahony2017geometric, van2019geometric, mahony2021homogeneous], and have led to applications in various fields, e.g.,  [barrau2018invariant, van2020invariant, hartley2020contact, pavlasek2021invariant, wu2017invariant, heo2018consistent, mahony2017geometric] and in the industry [barrau2018invariant]. The field of observers has benefited from the introduction of new groups brought by the IEKF theory, see e.g., [wang2020hybrid, hashim2021gps, van2021autonomous], namely the groups $SE_{2}(3)$ and $SE_{k}(d)$ introduced in [barrau2015non, barrau2016invariant] and their relation to the navigation equations, see [barrau2022geometry] for a recent perspective. The equivariant observer framework, see [mahony2021equivariant], is closely related.
@@
In this technical note, we introduce the Iterated Invariant Extended Kalman Filter (IterIEKF), a refinement of the IEKF [barrau2018invariant] that leverages the Gauss-Newton (GN) method to enhance its measurement update step, and which systematically improves its accuracy. Our main contribution is then a comprehensive analysis of the IterIEKF properties, revealing rare properties in nonlinear estimation that remind of the linear case, and offering new tools and insights into the theory of invariant filtering. Finally, to highlight the practical relevance of our approach, we apply the IterIEKF to a problem of engineering interest, and show it outperforms the IEKF when observations are very accurate (low measurement noise).
@@
While the ultimate goal of invariant filtering is to recover properties of the linear Kalman filter in a nonlinear setting, and to identify systems of interest that lend themselves to this goal, a parallel line of work has been concerned with how to accommodate in general the geometric nature of the state space, without seeking further properties. In particular, the “Lie-Group EKF” (LG-EKF) from [bourmaud2013discrete] proposes a generic intrinsic version of the EKF on Lie groups, that accounts for geometric features such as curvature. An iterated version, called LG-IterEKF was proposed in [bourmaud2016intrinsic]. See, e.g., [he2021kalman] and applications to inertial-lidar navigation [xu2022fast] for recent references in this field. When tailored to the specific class of problems we consider, namely by using the groups from invariant filtering and the choice of errors that it advocates, the LG-IterEKF (that we call then adapted) is close to the proposed IterIEKF, although slightly different. As a result, we can prove the properties of the IterIEKF carry over to this adapted LG-IterEKF. This novel result constitutes a secondary contribution.
@@
Section II introduces the proposed IterIEKF algorithm for both noisy and noise-free measurements. Section III derives theoretical properties of the IterIEKF in the context of noise-free measurements. We prove they carry over to the LG-IterEKF from [bourmaud2016intrinsic] under some conditions. Section IV evaluates in simulations the performance of the proposed algorithm compared to other state-of-the-art filters for estimating the extended pose (orientation, velocity, and position) of a crane hook, equipped with an Inertial Measurement Unit (IMU).
@@ II-A The IEKF equations
The invariant framework assumes the state $\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}$ is an element of a matrix Lie group $G\subset GL_{N}(\mathbb{R})$ having dimension $n$, where $GL_{N}(\mathbb{R})$ denotes the group of $N\times N$ invertible matrices. An example of such a state $\boldsymbol{\chi}$ is given below, in (15). The exponential map of $G$ is defined as $\exp_{G}(\cdot):=\exp_{m}(\mathcal{L}_{\mathfrak{g}}(\cdot))$, where $\exp_{m}(\cdot)$ is the matrix exponential and $\mathcal{L}_{\mathfrak{g}}(\cdot)$ is the bijective linear map identifying the Lie algebra $\mathfrak{g}$ with $\mathbb{R}^{n}$, see e.g., [barrau2016invariant]. The IEKF comes in two versions, the left- or the right-invariant version, depending on the form of the observations, see [barrau2018invariant]. This note focuses on the left-IEKF. Transposition to the right-IEKF is straightforward.
@@
The (left) IEKF assumes that the state follows a concentrated Gaussian distribution on $G$ [chirikjian2009stochastic, wang2006error, wolfe2011bayesian, bourmaud2015continuous]. At time $k$, letting $l$ denote current time index $k$, or previous index k1k-1,
@@
When the function $\mathbf{f}$ possesses the group affine property (i.e., is of the form $\mathbf{f}(\raisebox{1.0pt}{$\boldsymbol{\chi}$},\mathbf{u},\mathbf{w})=\bar{\mathbf{f}}(\raisebox{1.0pt}{$\boldsymbol{\chi}$},\mathbf{u})\mathbf{g}(\mathbf{w})$ where $\bar{\mathbf{f}}$ satisfies $\bar{\mathbf{f}}(\mathbf{a}\mathbf{b},\mathbf{u})=\bar{\mathbf{f}}(\mathbf{a},\mathbf{u})\bar{\mathbf{f}}(\mathbf{I},\mathbf{u})^{-1}\bar{\mathbf{f}}(\mathbf{b},\mathbf{u})$, for all $\mathbf{a},\mathbf{b}\in G$, $\mathbf{u}\in\mathbb{R}^{m}$, see [barrau2018invariant]), the Jacobian $\mathbf{F}_{k}$ becomes independent of the current state estimate $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}$ [barrau2016invariant, barrau2018invariant]. Moreover, in the absence of process noise, this results in the exact error dynamics $\boldsymbol{\xi}_{k+1\mid k}=\mathbf{F}_{k}\boldsymbol{\xi}_{k\mid k}$, where the Jacobian $\mathbf{F}_{k}$, i.e., the first order, fully captures the nonlinearity of $\mathbf{f}$. This property called log-linearity is a key feature of invariant filtering. Note that an IEKF can be devised even if $\mathbf{f}$ is not group affine, based on expansion (4), and the theory to follow applies.
@@ II-B The IterIEKF equations
The optimization problem (6) does not admit a closed-form solution. The IEKF addresses this by approximating $\exp_{G}(\boldsymbol{\xi})$ using its first-order Taylor expansion around $\mathbf{0}$. Inspired by the iterated EKF [bell1993iterated], a more accurate solution can be obtained by applying the GN algorithm to iteratively refine the estimate until convergence. This process results in a sequence of updates that closely resemble those of the IEKF. The resulting Iterated IEKF (IterIEKF) is outlined in Algorithm 1. Details are provided in Appendix A-A.
@@
$\mathbf{\mathcal{J}}_{r}(\boldsymbol{\xi})$ denotes the right Jacobian [chirikjian2009stochastic] of $G$ at $\boldsymbol{\xi}$, see (30).
Iterating proves particularly useful when measurement noise is low, as we will see in the sequel. However, in the limit case where the measurement noise magnitude approaches zero, the innovation covariance $\mathbf{S}_{k}$ may become rank-deficient [goffin2023invariant], in which case the Kalman gain is undefined. An adapted version of the IterIEKF for this specific scenario is detailed in Algorithm 2. Details are provided in Appendix A-B.
@@ III Theoretical properties of the IterIEKF
Deriving properties in the presence of measurement noise seems out of reach, as is often the case in nonlinear filtering (e.g., the log-linear error property of [barrau2016invariant] only holds in the absence of process noise). However, we show that the IterIEKF possesses strong properties in the absence of measurement noise (whereas the state estimate and the propagation may be noisy). We therefore consider noise-free measurement
@@ III-A Properties of considered noise-free measurements
The theory of (left) invariant filtering focuses on measurements of the form (1b), see [barrau2016invariant], which boil down to (7) when the noise is turned off. It turns out that the corresponding “observed set” (8) possesses an interesting structure.
@@
The first point of Proposition 2 is directly useful for proving the compatibility properties of the IterIEKF, in the sense of Definition 2. This will be established in the next subsection, relying on the following key Lemma 1. We also note, in passing, that the second point will only become relevant when extending the present theory to the iterated EKF on Lie groups (LG-IterEKF) of [bourmaud2015continuous], under certain conditions.
@@ III-C Consequences of the compatibility property
Let us first consider the first point of Definition 2. During the update with $\mathbf{y}_{k}$, we see at line 17 of Algorithm 1 that $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}=\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k-1}\exp_{G}(\boldsymbol{\xi}_{k\mid k-1}^{i})$, where $\boldsymbol{\xi}_{k\mid k-1}^{i}=\mathbf{K}_{k}^{i}\mathbf{z}_{k}^{i}\in\operatorname{Im}\mathbf{K}_{k}^{i}\subseteq\operatorname{Im}\mathbf{P}_{k\mid k-1}$, with $\operatorname{Im}\mathbf{K}_{k}^{i}$ denoting the span of $\mathbf{K}_{k}^{i}$. Since $\bar{\mathbf{H}}_{k}\mathbf{P}_{k\mid k-1}\bar{\mathbf{H}}_{k}^{T}=\mathbf{0}$ by assumption, we necessarily have $\bar{\mathbf{H}}_{k}\boldsymbol{\xi}_{k\mid k-1}^{i}=\mathbf{0}$. Then, as $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k-1}\in\mathcal{S}_{\{\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}\bar{\mathbf{d}}_{k}=\bar{\mathbf{y}}_{k}\}}$, a direct application of Lemma 1 ensures $\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k\mid k}\in\mathcal{S}_{\{\raisebox{1.0pt}{$\boldsymbol{\chi}$}\bar{\mathbf{d}}_{k}=\bar{\mathbf{y}}_{k}\}}$.
@@ III-D Extending the theory to the iterated EKF on Lie groups
The LG-IterEKF, introduced in [bourmaud2016intrinsic], transposes the standard iterated EKF to Lie group state spaces, by leveraging an intrinsic GN method on the group, that accounts for its manifold structure, notably curvature. LG-IterEKF of [bourmaud2016intrinsic] is a general algorithm, meant to address general problems on Lie groups having no specific structure, beyond the intrinsic nature of the state space. By contrast, our IterIEKF addresses specific types of observations, namely (1b) (or its right-invariant counterpart $\mathbf{y}_{k}=\raisebox{1.0pt}{$\boldsymbol{\chi}$}_{k}^{-1}\mathbf{d}_{k}+\mathbf{n}_{k}$, in which case an iterated right-invariant EKF should be used, by mimicking the difference between the LIEKF and the RIEKF of [barrau2016invariant]). It defines a modified innovation $\mathbf{z}_{k}=\hat{\raisebox{1.0pt}{$\boldsymbol{\chi}$}}_{k}^{-1}\mathbf{y}_{k}-\mathbf{d}_{k}$, allowing for the use of the invariant Kalman filter framework of [barrau2016invariant, barrau2018invariant]. The standard invariant EKF has had various successes across control, navigation, and robotics, and is clearly a different algorithm than the LG-EKF of [bourmaud2013discrete].
That said, the questions one could address are as follows. Assume we use the theory of invariant filtering to endow the LG-IterEKF with all the features of the IterIEKF—namely, measurements of the form (7), uncertainty model (2), and the Kalman gain from Algorithm 2 to address rank deficiency issues, as well as the appropriate choice regarding left or right versions (note that [bourmaud2016intrinsic] chooses a systematic right-invariant error based setting whereas we advocate to choose the error depending on the form of the observations), resulting in an “invariant” version of the LG-IterEKF, that we call adapted LG-IterEKF, see Algorithm 3. Then, 1) do we actually recover the IterIEKF, and 2) if no, does it inherit any of the properties we have proved, given that no such properties have ever been proved for the LG-IterEKF? The answer to point 1 is no, as intrinsic GN on the group entails adding various Jacobians and differs from the GN algorithm addressing (6), that directly optimizes in a linear space. The answer to point 2 is yes, the properties above do carry over. This is detailed, proved, and discussed, in Appendix B.
Algorithm 3 The adapted LG-IterEKF from [bourmaud2016intrinsic] when measurements (7) are considered, uncertainty model (2) is used, potential rank deficiency issues in the Kalman gain are addressed via the noise-free gain from Algorithm 2, and a left-invariant error is used instead. See (30) for definition of $\mathbf{\mathcal{J}}_{r}$.